OpenStudy (anonymous):

How do I solve the log equation? 3 log 2x= 4

OpenStudy (tkhunny):

Standard operations. Several ways to go about it. 1) Divide by 3 2) Transform to Exponential Form. 3) Divide by 2 OR 1) Use log rules to move '3' inside the logarithm 2) Transform to Exponential Form. 3) Find the cube root. 4) Divide by 2 Pick a path that seem good to you,

OpenStudy (anonymous):

log 2x=1.33.....log2+logx=1.33 hence find x..

OpenStudy (tkhunny):

Do we want a numerical answer with ambiguous rounding? \[x = \frac{1}{2}\cdot 10^{\frac{4}{3}}\] This is an exact answer.

OpenStudy (anonymous):

Icame up with x=10.77 right?

OpenStudy (tkhunny):

10.772173450159418608796467832596 is closer. :-) Good work. Let's see another one on your own.